# Suppressing Overestimation in Q-Learning through Adversarial Behaviors

### v1

@@ A.3 Weng’s Example
Figure 10 shows the experimental results using Weng’s Example. In Figure 9(a), we give results with varying shifts. We tested with a scale from $10^{1}$ to $10^{-4}$, both for positive and negative values of shifts. As the value of shifts gets closer to zero, DAQs converge faster. The result with constant step-size is shown in Figure 9(b). The graphs are very similar to Figure 4 except for the behavior of Q-learning.
@@
(a) Results with various shift values, which are presented in titles of each subfigures. $\alpha_{n}=\frac{10}{n+100}$ and $M=8$ is used. The graphs of Q-learning, Double Q-learning, minmax Q-learning and maxmin Q-learning are the same for all six subfigures for better comparison. From upper left to bottom right, the amount of shifts decreases and the performance of DAQs gets better.
@@
(b) Result using constant step-size $\alpha=0.1$. From the first subfigure to the last, $M=8,20$ and $100$, respectively. $b_{1}=+0.001$ and $b_{2}=+0.002$ are used for both DAQs.
Figure 10: Result using Weng’s Example. A 5% line is optimal. $\epsilon=0.1$ and $\gamma=1$ are used. Data is averaged over 10,000 experiments.

### v3

@@ A-C Weng’s MDP
Fig. 15 and Fig. 16 show the experimental results using Weng’s MDP. A 5% line is optimal. In Fig. 15, we give results with varying shifts. We tested with a scale from $10^{1}$ to $10^{-4}$, both for positive and negative values of shifts. As the value of shifts gets closer to zero, DAQs converge faster. The result with constant step-size is shown in Fig. 16. The graphs are very similar to Fig. 6 except for the behavior of Q-learning.
@@
Figure 15: Results with various shift values, which are presented in titles of each subfigure. $\alpha_{n}=\frac{10}{n+100}$ and $M=8$ is used. The graphs of Q-learning, double Q-learning, minmax Q-learning and maxmin Q-learning are the same for all six subfigures for better comparison. From upper left to bottom right, the amount of shifts decreases and the performance of DAQs gets better.
@@
Figure 16: Result using constant step-size $\alpha=0.1$. From the first subfigure to the last, $M=8,20$ and $100$, respectively. $b_{1}=+0.001$ and $b_{2}=+0.002$ are used for both DAQs.